OpenStudy (anonymous):

8x + 2y = -4 It says to graph the equation. I tried to find help online but it seems to need to be in a different order. This is actually for a final study guide. I used to know how to do it, so i need a refresher.

OpenStudy (anonymous):

First what you have to do is arrange the equation so you get the basic y=mx+b order. To do this, subtract 8x from both sides of the equation. Now you have 2y=-8x-4. To get y, we need to divide by 2 on both sides. Now we have y=-4x-2, and that's the equation you'll graph. To graph it, you take your x coordinate, for example 1, multiply it by -4 and then subtract 2, then that will be the y coordinate. I hope you know how to graph the equation or that my summary is good enough.

OpenStudy (anonymous):

Thanks, you helped a lot. Ill try it out. I feel so awful, i dont remember most of this math crap for my final which is tomorrow XD Have a good day

OpenStudy (anonymous):

Put the equation into slope intercept form. y= mx + b m being slope b being the y-intercept 8x + 2y = -4 divide by 2y on both sides 8x/ 2 -4/2 4x-y=-2 add y to both sides put in slope intercept form your final answer would be y=-4x-2